In the electric field of a point charge ‘q’ ,the four points A,B,C and D are equidistant from q, however AB>AC>AD.Calculate the work donr in taking a unit charge along AB,AC and AD.

Grade:12

1 Answers

Vikas TU
6 years ago
Dear Student,
As all the points are equidistant.
Work done at point B = kq/r
Work done at point A =kq/r
Work done at point C = kq/r
Work done at point D = kq/r.
Cheers!!
Regards,
Vikas (B. Tech. 4th year
Thapar University)